### v1

# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

Abstract
Text-guided diffusion models catalyze a paradigm shift in audio generation, facilitating the adaptability of source audio to conform to specific textual prompts. Recent advancements introduce inversion techniques, like DDIM inversion, to zero-shot editing, exploiting pre-trained diffusion models for audio modification. Nonetheless, our investigation exposes that DDIM inversion suffers from an accumulation of errors across each diffusion step, undermining its efficacy. And the lack of attention control hinders the fine-grained manipulations of music. To counteract these limitations, we introduce the Disentangled Inversion technique, which is designed to disentangle the diffusion process into triple branches, thereby magnifying their individual capabilities for both precise editing and preservation. Furthermore, we propose the Harmonized Attention Control framework, which unifies the mutual self-attention and cross-attention with an additional Harmonic Branch to achieve the desired composition and structural information in the target music. Collectively, these innovations comprise the Disentangled Inversion Control (DIC) framework, enabling accurate music editing whilst safeguarding structural integrity. To benchmark audio editing efficacy, we introduce ZoME-Bench, a comprehensive music editing benchmark hosting 1,100 samples spread across 10 distinct editing categories, which facilitates both zero-shot and instruction-based music editing tasks. Our method demonstrates unparalleled performance in edit fidelity and essential content preservation, outperforming contemporary state-of-the-art inversion techniques. [^1] Audio samples are available at https://MEDIC-Zero.github.io/. [^2] Benchmark data are available at https://huggingface.co/datasets/liuhuadai/ZoME-Bench.
1 Introduction
Advancements in text-guided diffusion models (Song et al., 2020; Song & Ermon, 2020; Peebles & Xie, 2023) have markedly enhanced audio generation, leveraging their impressive capability for realistic and varied outputs. These models (Liu et al., 2023b; Huang et al., 2023; Liu et al., 2024) provide the foundation for prompt-based audio editing, offering new opportunities to modify audio landscapes in response to specific textual prompts. Early audio editing strategies rely on training models from scratch or test-time optimization, both of which are hampered by intensive computational demands. Recent advances (Manor & Michaeli, 2024; Zhang et al., 2024) in zero-shot audio editing have been made through DDIM (Song et al., 2020) and DDPM (Ho et al., ) inversion techniques, but challenges remain.
Key among these challenges is maintaining editing fidelity - ensuring that the model’s output matches the given instructions - and preserving essential content, a hallmark of effective audio editing that targets only necessary changes. Balancing these objectives involves a careful exchange of information between the source and target branch in diffusion processes, with existing inversion methods like DDIM proving suboptimal for conditional diffusions (Mokady et al., 2023). Enhanced versions of edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024) make strides in preservation by imprinting the source onto the noise space, yet at the expense of reduced modification capabilities due to noise reduction.
In this work, we rigorously examine the shortcomings of the DDIM inversion approach. Our comprehensive analysis, suggests that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence. The implementation of Classifier-free Guidance (CFG) (Ho & Salimans, 2021) further amplifies the accumulated errors. Moreover, attention control (Cao et al., 2023; Hertz et al., 2022) has demonstrated its power in achieving high fidelity and essential content preservation. Yet, analogous control mechanisms in music editing are noticeably absent. To bridge this gap, we introduce an innovative Disentangled Inversion Control technique. This technique is comprised of two main components: Harmonic Attention Control and Disentangled Inversion. The former integrates cross-attention and mutual self-attention controls, supplemented by a novel Harmonic Branch, which facilitates precise, fine-grained manipulations for both rigid and non-rigid editing scenarios. The latter component strategically deconstructs the diffusion process into three distinct branches, enabling each to realize its full potential independently.
Acknowledging the lack of standardized benchmarks in the field of audio editing, we introduce ZoME-Bench. The first music editing benchmark consists of 1,100 audio clips, distributing them into 10 rigorously curated editing categories. Each entry is carefully assembled, comprising a source prompt, a target prompt, human instruction, and blended words intended for editing. Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
Our contributions can be summarized as follows. 1) We introduce a novel, tuning-free methodology referred to as Disentangled Inversion Control (DIC), designed to facilitate both consistent manipulations of musical elements and intricate non-rigid editing tasks. 2) An Harmonic Attention Control framework is introduced to unify cross-attention control and self-attention control, which enables both rigid and non-rigid editing. 3) We present Disentangled Inversion Technique to achieve superior results with negligible inversion error without optimization by branch disentanglement, aiding in accurately editing the music while preserving the structural information. 4) We introduce a tailored benchmark for music editing, named ZoME-Bench, and Experimental results on ZoME-Bench demonstrate that DIC outperforms baselines, achieving substantial gains in both essential content preservation and editing fidelity.
2 Related Works
2.1 Text-based Audio editing
The objective of text-based audio editing (Paissan et al., 2023; Plitsis et al., 2024; Han et al., 2023a) is to utilize diffusion models to manipulate audio content based on the provided target prompt. Two primary challenges command intense focus. The first is the fidelity of editing, which pertains to the model’s ability to strictly follow the initial instructions. The second challenge involves essential content preservation, a critical attribute of competent audio editing models that ensures only suitable regional modifications are made during the editing process. Existing methodologies for addressing these intricate challenges typically follow one of three paths. The first involves attempts to develop end-to-end editing models (Copet et al., 2023; Agostinelli et al., 2023; Chen et al., 2024)that employ diffusion processes. However, these efforts are often hampered by indirect training strategies or a lack of comprehensive datasets. The second path involves test-time optimization strategies that utilize large pre-trained models for editing (Paissan et al., 2023; Plitsis et al., 2024). Despite their versatility, these methods are often burdened by the significant computational demands of fine-tuning diffusion models or optimizing text-embeddings for signal reconstruction. Some methods may choose to employ both strategies (Kawar et al., 2023), further increasing the computational load. The final path involves inversion techniques, which typically use DDPM (Huberman-Spiegelglas et al., 2024; Wu & De la Torre, 2023)/DDIM (Song et al., 2020) inversion strategies to extract diffusion noise vectors that match the source signal. Given its rapid and intuitive zero-shot editing capabilities, we have chosen inversion techniques as our primary research framework. In this study, we propose a new inversion technique named Disentangled Inversion Control. This technique aims to achieve accurate editing while preserving structural information.
2.2 Inversion Techniques
The field of image inversion techniques has experienced significant progress in recent years (Brooks et al., 2023; Kim et al., 2022; Parmar et al., 2023; Dhariwal & Nichol, 2021). While the DDIM inversion proves effective for unconditional diffusion models (Song et al., 2020; Wallace et al., 2023) , its limitations become apparent when applied to text-guided diffusion models, particularly when classifier-free guidance is necessary for meaningful editing. A range of solutions (Mokady et al., 2023; Tumanyan et al., 2022) have been proposed to address these challenges. For example, Negative-Prompt Inversion strategically assigns conditioned text embeddings to Null-Text embeddings, effectively reducing potential deviation during the editing process. Conversely, Edit Friendly DDPM provides an alternative latent noise space via modified DDPM sample distributions, promoting the successful reconstruction of the desired image (Huberman-Spiegelglas et al., 2024).
Recently, optimization-based inversion methods using specific latent variables have gained popularity (Ju et al., 2023; Kawar et al., 2023). These are designed to minimize accumulated errors stemming from the DDIM inversion. Techniques such as Null-Text Inversion (Mokady et al., 2023) are promising, but they introduce complexity and instability into the optimization process, making it relatively time-consuming. Differently, we introduce a plug-in-plus method called Disentangled Inversion Control to separate branches, achieving superior performance with considerably fewer computational resources.
Figure 1: The framework and rationale behind Disentangled Inversion Control. Traditional DDIM inversion introduces incremental errors at each step, leading to only a coarse approximation of the source audio’s distribution. The implementation of classifier-free guidance exacerbates the error accumulation. In contrast, our method intricately separates the diffusion process into a triple-branch system, allowing each branch to fully realize its individual potential.
3 Disentangled Inversion Control
3.1 Motivation
Figure 1 and Preliminaries in Appendix 2 reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising the original’s essence. The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
In the landscape of prompt-based editing (Dong et al., 2023; Kim et al., 2022; Feng et al., 2023), the ability to grasp the subtleties of linguistics and enable more granular cross-modal interactions stands as a formidable challenge.  Hertz et al. (2022) acknowledges that in image editing, the fusion between text and visual modalities happens within the parameterized noise prediction network $\epsilon_{\theta}$. This sparks a plethora of attention control techniques that direct the image domain’s target denoiser network $\hat{\epsilon}_{\theta}$ to better reflect target prompts. Yet, analogous control mechanisms in music editing are noticeably absent.
Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach poised to achieve fine-grained zero-shot music editing. DIC strategically disentangles the diffusion process as a triple-branch, granting each branch to optimize its functionality. At the same time, the strategy leverages harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit relevance. We will first introduce Harmonized Attention Control in Section 3.2 and discuss Disentangled Inversion in Section 3.3.
3.2 Harmonized Attention Control Framework
The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023a). At the denoising step $t$, the output of the $(l-1)$-th block is passed through self-attention and then aligned with textual cues from prompt $P$ within the cross-attention layer. The attention mechanism is formalized as:
$\text{Attention}(Q,K,V)=MV=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (1)
where $Q$ denotes the query features derived from the audio features, $K$, $V$ represent the key and value features, and $M$ is the attention maps. We explore varying semantic transformations of audio content through bifurcated attention control strategies—cross-attention for rigid changes and mutual self-attention for non-rigid adjustments. Finally, we introduce an intermediate branch to host the desired composition and structural information in the target music.
Figure 2: The framework of Harmonized attention control (HAC). HAC unifies cross-attention control and mutual self-attention control with an additional branch named Harmonic Branch to host the desired composition and structural information in the target music.
3.2.1 Cross-Attention Control
Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt ${\mathcal{P}}$, into a second generation with the modified prompt ${\mathcal{P}}^{*}$. We follow the practice of Prompt-to-Prompt and define CAC as Global Attention Refinement and Local Attention Blend.
Global Attention Refinement
At a given time step $t$, the attention map $M_{t}$ for both the origin and target branches is computed, averaging over all layers with respect to the noised latent $\mathbf{z}_{t}$. We employ an alignment function $A$ that maps each token index from the target prompt $\mathcal{P}^{*}$ to its equivalent in $\mathcal{P}$ or to None for non-aligning tokens. The refinement action is thus:
$\text{Refine}(M^{src}_{t},M^{tgt}_{t},t)=\begin{cases}(M^{tgt}_{t})_{i,j}&\text{if }A(j)=\text{None},\\ (M_{t})_{i,j}&\text{otherwise}.\end{cases}$ (2)
Local Attention Blends
Beyond global attention, we incorporate a blending mechanism as suggested by  Hertz et al. (2022) and  Mokady et al. (2023). This method selectively integrates and maintains certain semantics by using target blend words $w^{tgt}$ for semantic additions and source blend words $w^{src}$ for semantic preservation. At each denoising step $t$, the mechanism operates on the target latent ${\bm{z}}^{tgt}_{t}$ as follows:
$m_{\text{tgt}}$ $=\text{Threshold}\left[M^{tgt}_{t},w_{\text{tgt}},k_{\text{tgt}}\right],$ (3)
$m_{\text{src}}$ $=\text{Threshold}\left[M^{src}_{t},w_{\text{src}},k_{\text{src}}\right],$ (4)
${\bm{z}}^{tgt}_{t}$ $=(1-m^{tgt}+m^{src})\odot{\bm{z}}^{src}_{t}+(m^{tgt}-m^{src})\odot{\bm{z}}^{tgt}_{t}$ (5)
where $m_{\text{tgt}}$ and $m_{\text{src}}$ are binary masks and threshold function is as delineated below:
$\text{Threshold}(M,k)=\begin{cases}1&\text{if }M_{i,j}\geq k,\\ 0&\text{if }M_{i,j}<k.\end{cases}$ (6)
For simplificity, we define the process of local editing in Equation 10 as:
${\bm{z}}^{tgt}_{t}=\text{LocalEdit}({\bm{z}}^{src}_{t},{\bm{z}}^{tgt}_{t},M^{src}_{t},M^{tgt}_{t},w_{\text{src}},w_{\text{tgt}})$ (7)
Scheduling Cross-Attention Control
Applying cross-attention control only at early stages ensures creative flexibility while maintaining structure. Acting on insights from  Hertz et al. (2022), we limit cross-attention to the initial phases up to a cutoff point $\tau_{c}$. This moderation allows us to capture the nuances and intended changes in musical compositions effectively. The approach is defined as:
$\text{CrossEdit}(M_{t},M^{*}_{t},t)=\begin{cases}\text{Refine}(M^{src},M^{tgt})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
3.2.2 Mutual Self-Attention Control
We diverge from the conventional use of cross-attention mechanisms and instead draw inspiration from the MasaCtrl (Cao et al., 2023) technique to refine music structure through self-attention queries. These queries adeptly navigate through non-rigid musical transformations, aligning with the designated musical theme or instrument (target prompt). The process begins by sketching the foundational musical theme using the target’s self-attention components—$Q^{tgt}$, $K^{tgt}$, and $V^{tgt}$. This is followed by enriching this theme with elements resembling the thematic content from the source ($K^{src}$, $V^{src}$), steered by $Q^{tgt}$. However, applying this attentive modulation uniformly over all processing layers and through every denoising step might result in a composition excessively mirroring the source. Consequently, echoing the ethos of MasaCtrl, our proposed solution selectively employs mutual self-attention in the decoder portion of our music editing U-Net, initiated after a set number of denoising iterations.
Scheduling Mutual Self-Attention Control
The application of mutual self-attention is meticulously planned, beginning at a specific denoising step $S$ and extending beyond a designated layer $L$. The strength and influence of this control mechanism are designed as follows:
$\text{SelfEdit}({Q^{src},K^{src},V^{src}},{Q^{tgt},K^{tgt},V^{tgt}},t)=\begin{cases}{Q^{src},K^{src},V^{src}}&\text{if }t\geq S\text{ and }l\geq L,\\ {Q^{tgt},K^{src},V^{src}}&\text{otherwise}\\ \end{cases}$ (9)
In this framework, $S$ signifies the denoising step from which the mutual self-attention control commences, serving as a temporal threshold. Similarly, $L$ distinguishes the layer index below which this nuanced control strategy becomes operational, tailoring the musical output towards the intended artistic direction.
3.2.3 Harmonic Branch Integretion
The naive combination of cross-attention control and mutual self-attention control sequentially would lead to a sub-optimal outcome in the original dual-branch setup, especially failing the global attention refinement. To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate to host the desired composition and structural information in the target music. Our unified framework is detailed in Algorithm 2 and illustrated in Figure 2. During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$. Following this, cross-attention control is applied on $M_{har}$ and $M_{tgt}$ to refine the semantic information for $M_{tgt}$.
Algorithm 1 Harmonized Attention Control in one DDIM Forward Process
1:  Input: A source prompt ${\mathcal{P}}$, a target prompt ${\mathcal{P}}^{*}$, a source audio latent ${\bm{z}}_{0}$, denoising network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$, current time step $\tau$, source and target blend words $w_{\text{src}},w_{\text{tgt}}$, input latents $z^{\text{src}}_{\tau},z^{\text{tgt}}_{\tau},z^{\text{har}}_{\tau}$.
2:  $\epsilon_{\text{src}},\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},M_{\text{src}}=\epsilon_{\theta}(z^{\text{src}}_{\tau},\tau,c_{\text{src}})$
3:  $\epsilon_{\text{tgt}},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},M_{\text{tgt}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}})$
4:  $\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\}=\text{SelfEdit}(\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},\tau)$
5:  $\epsilon_{\text{har}},M_{\text{har}}=\epsilon_{\theta}(z^{\text{har}}_{\tau},\tau,c_{\text{src}};\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\})$
6:  $\hat{M^{\text{tgt}}}=\text{CrossEdit}(M_{\text{har}},M_{\text{tgt}},\tau)$
7:  $\hat{\epsilon_{\text{tgt}}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}};\hat{M^{\text{tgt}}})$
8:  $z^{\text{src}}_{\tau-1},z^{\text{tgt}}_{\tau-1},z^{\text{har}}_{\tau-1}=\text{Sample}([z^{\text{src}}_{\tau},z^{\text{tgt}}_{\tau},z^{\text{har}}_{\tau}],[\epsilon_{\text{src}},\epsilon_{c_{\text{tgt}}},\epsilon_{\text{har}}],\tau)$
9:  ${\bm{z}}^{tgt}_{\tau-1}$ = $\text{LocalEdit}({\bm{z}}^{src}_{\tau-1},{\bm{z}}^{tgt}_{\tau-1},M^{src}_{\tau-1},M^{tgt}_{t},w_{src},w_{tgt})$
10:  Output: $z^{\text{src}}_{\tau-1},z^{\text{tgt}}_{\tau-1},z^{\text{har}}_{\tau-1}$
3.3 Disentangled Inversion Technique
Recognizing the limitations of using DDIM inversion without classifier-free guidance, we observe that it yields an easily modifiable but imprecise approximation of the original audio signal. Increasing the guidance scale enhances editability but sacrifices reconstruction accuracy due to latent code deviation during editing.
In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the target, and the harmonized branch. This bifurcation is designed to optimize the performance of each branch separately. For the source branch, we implement a targeted correction mechanism. By reintegrating the distance $z^{*}_{t}-z^{src}_{t}$ into $z^{src}_{t}$, we directly mitigate the deviation of the pathway. This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio. On the other hand, the target branch and harmonic branch is left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio. The branches’ untouched state ensures that the model’s potential is utilized to its fullest extent, thereby ensuring the fidelity and integrity of the generated audio. We will further discuss this in the Section 5.4.2. The algorithm of Disentangled Inversion Technique has been outlined in Algorithm 2.
Typical diffusion-based editing (Han et al., 2023b; Miyake et al., 2023) involves two parts: an inversion process to get the diffusion space of the audio, and a forward process to perform editing on the diffusion space. Disentangled Inversion can be plug-and-played into the forward process and rectifies the deviation path step by step. Specifically, Disentangled Inversion first computes the difference between ${\bm{z}}^{*}_{t-1}$ and ${\bm{z}}^{src}_{t-1}$, then adds the difference back to ${\bm{z}}^{src}_{t-1}$ in DDIM forward. We only add the difference of the source prompt in latent space and update ${\bm{z}}^{src}_{t-1}$, which is the key to retaining the editability of the target prompt’s latent space.
Algorithm 2 Disentangled Inversion Technique
1:  Input: A source prompt ${\mathcal{P}}$, a target prompt ${\mathcal{P}}^{*}$, a source audio latent ${\bm{z}}_{0}$, and guidance scale $\omega$.
2:  Output: A edited audio latent ${\bm{z}}^{tgt}_{0}$.
3:  Compute the intermediate results ${\bm{z}}^{*}_{T},...,{\bm{z}}^{*}_{1}$ using DDIM inversion over ${\bm{z}}_{0}$.
4:  Initialize ${\bm{z}}^{src}_{T}\leftarrow{\bm{z}}^{*}_{T}$, ${\bm{z}}^{tgt}_{T}\leftarrow{\bm{z}}^{*}_{T}$, ${\bm{z}}^{har}_{T}\leftarrow{\bm{z}}^{*}_{T}$.
5:  for $t=T$ to $1$ do
6:     $[{\bm{d}}^{src}_{t-1},{\bm{d}}^{tgt}_{t-1},{\bm{d}}^{har}_{t-1}]\leftarrow{\bm{z}}^{*}_{t-1}$ DDIM_Forward$({\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega)$
7:     ${\bm{z}}^{src}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
8:     ${\bm{z}}^{har}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{har}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
9:     ${\bm{z}}^{tgt}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{tgt}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
10:  end for
11:  return  ${\bm{z}}^{tgt}_{0}$
4 BenchMark Construction
Despite recent significant advancements in music generation utilizing the latent diffusion model, music editing has received comparably less attention, especially zero-shot music editing. Hence, to systematically validate our proposed method as a plug-and-play strategy for editing models and compare our method with existing zero-shot music editing methods, as well as compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset, named ZoME-Bench (Zero-shot Music Editing Benchmark). ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps, spanning ten distinct editing types such as add instrument, delete instrument, change genre, and so on. Each sample in this collection is accompanied by its corresponding source prompt, target prompt, human instruction, and the source audio. Additionally, annotations pertinent to attention control methodologies, such as blended words, are meticulously included to facilitate detailed evaluations. We attach the details about our benchmark in Appendix B.
5 Experiments
5.1 Implementation Details
We perform the inference of different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023b) models with 200 inference steps. The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA 4090Ti GPU for a fair comparison. We evaluate all methods in ZoMo-Bench across all editing types. Details can be found in Appendix C.
5.2 Evaluation Metrics
In order to evaluate the performance of our novel Disentangled Inversion Control framework, we employ a diverse set of five metrics that encompass three critical dimensions of audio quality and fidelity: Struture Distance,CLAP Score,LPAPS,MSE (Mean Squared Error)and FAD (Fréchet Audio Distance). All the metrics are introduced in detail in Appendix C.1.
5.3 Zero-shot Music Editing Results
[…]
SDEdit 25.87 0.22 12.18 9.82 0.40
DDPM-Friendly 22.36 0.21 14.08 8.49 0.41
DDIM Inversion 17.11 0.18 6.24 6.12 0.55
DIC 11.97 0.15 2.49 4.54 0.60
Table 1: Comparison with baselines on ZoMo-Bench about struture and content preservation and CLAP similiarity. DIC: our proposed Disentangled Inversion Control.
We present a comparative study of our Disentangled Inversion Control(DIC) against several established music editing techniques, namely SDEdit (Liu et al., ), DDIM Inversion Ho et al. , and DDPM-Friendly (Manor & Michaeli, 2024). Utilizing the ZoMo-Bench test set we released, we assess the quality of the generated audio samples. The assessment primarily concentrates on two key aspects: Essential Content Preservation and Editing Fidelity. The results are summarized in Table 1, leading to two main conclusions:
(1) Our Disentangled Inversion Control demonstrates substantial advancements in maintaining essential content and enhancing editing fidelity, in comparison to the baseline methods. (2) Despite DDIM Inversion achieving high CLAP similarity scores, significantly underperforms in preserving essential content when compared to our approach.
(a) Spider chart for editing quality.
(b) Spider chart for editing consistency.
Figure 3: A comprehensive performance evaluation on the ZoME-bench. We present spider charts of editing quality (CLAP Scores) and consistency (LPAPS) across 9 editing tasks for DDPM-Friendly, DDIM Inversion, and Disentangled Inversion Control methods.
To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing. The LPAPS and CLAP scores are depicted in Figure 3. The insights gleaned from this analysis are as follows: (1) Across all editing types, our Disentangled Inversion Control surpasses other methods, demonstrating its prowess in handling both rigid and non-rigid editing tasks. (1) While methods like DDIM Inversion exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing fine-grained manipulations, as reflected in their inferior performance, especially in the "Delete Instrument" and "Change Melody" non-rigid editing type.
5.4 Ablation Study
The ablation studies presented in this section aim to validate the contributions of the Unified Attention Control and Refining Dual-path Inversion to our framework’s overall performance. Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in the Appendix D.
5.4.1 Ablation on Attention Control Methods
[…]
HAC 11.97 0.15 2.49 4.54 0.60
w/ CAC 10.24 0.19 2.75 4.76 0.56
w/ MSAC 9.92 0.18 2.50 4.66 0.58
w/ HB 11.75 0.16 2.59 4.67 0.59
Table 2: Ablation results about different control methods. HAC: Harmonized Attention Control.
To ascertain the impact of our attention control mechanisms, we perform ablations on the following configurations: Cross-attention Control Only (w/ CAC), Mutual Self-attention Control only (w/ MSAC), without Harmonic Branch (w/o HB), and any attention control (w/o Control). The findings, detailed in Table 2, reveal that: (1) Both cross-attention and mutual self-attention controls individually enhance editing performance. (2) The combined use of CAC and MSAC further improves results, but the absence of the harmonic branch yields suboptimal outcomes.
5.4.2 Ablation on Disentangled Inversion Technique
To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from the Table 3: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score. 2) The incorporation of target distance into target latent and harmonic branch leads to an improvement in structure and content integrity but the CLAP score sharply decreases. 3) Add the harmonic distance into the harmonic latent witness a distant decrease.
[…]
$[{\bm{d}}_{src},{\bm{d}}_{src},{\bm{0}}]$ 14.28 0.17 2.64 4.67 0.57
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{src}]$ 13.17 0.17 2.55 4.59 0.58
$[{\bm{d}}_{src},{\bm{d}}_{tgt},{\bm{0}}]$ 11.24 0.15 2.44 4.32 0.56
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{tgt}]$ 11.12 0.14 2.41 4.29 0.56
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{har}]$ 37.51 0.27 27.2 14.23 0.28
$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.60
Table 3: Ablation results about disentangled inversion technique. [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance to the corresponding branch in Algorithm 2.
5.5 Qualitative Results
Figure 4: Visualizations of the source audio’s mel and edited mel-spectrograms by different editing methods.
To complement our quantitative findings, we present a qualitative comparison in Figure 4. Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with the preservation of melodic content and harmonic structure. In contrast, our Disentangled Inversion Control demonstrates superior performance in precise music editing while preserving structural integrity. We provide additional qualitative results in the Appendix G for further examination.
6 Conclusion
In this paper, we explored the burgeoning realm of text-guided diffusion models for audio generation, recognizing their potential to reshape source audio in alignment with specific textual prompts. Our retrospective analysis revealed that while recent methods like DDIM inversion have propelled zero-shot editing forward, they are not without their shortcomings. We identified that DDIM inversion is plagued by an accumulation of errors throughout the diffusion steps, which detracts from its overall effectiveness. Moreover, the absence of a fine-tuned attention control mechanism was found to impede the nuanced editing of music. To address these issues, we introduced the Disentangled Inversion framework, a novel approach that meticulously separates the source and target branches to amplify their individual strengths in preservation and editing. In tandem, we presented the Harmonized Attention Control technique, which enhances the preservation of essential content by integrating the attention maps from the original audio into the diffusion process. Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,000 samples across 10 varied editing categories—attested to the superiority of our methods.
Appendix A Preliminaries And Analyses
This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied in diffusion models for text-guided audio synthesis. It further explores the challenges associated with these methods.
A.1 Diffusion Models
Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, which corresponds to the given conditioning prompt. We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following the objective:
[…]
Where noise is added to the sampled data ${\bm{z}}_{0}$ according to timestamp $t$. At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using pre-trained network for $T$ steps. To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM samling:
${\bm{z}}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}{\bm{z}}_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ (11)
A.2 DDIM Inversion
While diffusion models have superior characteristics in the feature space that can support various down-stream tasks, it is hard to apply them to audios in the absence of natural diffusion feature space for non-generated audios. Thus, a simple inversion technique known as DDIM inversion is commonly used for unconditional diffusion models, predicated on the presumption that the ODE process can be reversed in the limit of infinitesimally small steps:
${\bm{z}}^{*}_{t}=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}{\bm{z}}^{*}_{t-1}+\sqrt{\alpha_{t}}(\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1})\epsilon_{\theta}({\bm{z}}^{*}_{t-1},t-1)$ (12)
However, in most text-based diffusion models, this presumption cannot be guaranteed, resulting in a perturbation from ${\bm{z}}_{t}$ to ${\bm{z}}^{*}_{t}$. Consequently, an additional perturbation from ${\bm{z}}^{*}_{t}$ to ${\bm{z}}^{src}_{t}$ arises when sampling an audio from ${\bm{z}}^{*}_{T}$ where $\alpha$ is hyper-parameter:
${\bm{z}}_{t}=\sqrt{\alpha_{t}}{\bm{z}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$ (13)
A.3 Classifier-free Guidance
Classifier-free Guidance (CFG) (Ho & Salimans, 2021) is proposed to overcome the limitation of text-conditioned models, where text adherence could be weak. However, a higher guidance scale, $\omega$, which is intended to strengthen the model’s fidelity to the text prompt, inadvertently magnifies the accumulated inversion error. This becomes problematic in editing scenarios where precise control over the audio synthesis is desired. The modified noise estimation in CFG can be expressed as:
$\hat{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}},\varnothing)=\omega\cdot\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})+(1-\omega)\cdot\epsilon_{\theta}({\bm{z}}_{t},t,\varnothing)$ (14)
where $\varnothing$ = ("") is the embedding of a null text. This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{"}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
Appendix B Benchmark construction
B.1 General information
Here are details about our ZoME-Bench dataset (Zero-shot Music Editing Benchmark).This dataset contains 1,000 audio samples , which are selected from MusicCaps , and each is 10 seconds long with a sample rate of 16k.
We refactored their captions to express an edit and divide them into 10 parts , each representing a type of editing. Sample and details are shown in the following table
Editing type id Editing type size origin prompt editing prompt editing instruction
0 change instrument 131 ambient acoustic [guitar] music ambient acoustic [violin] music change the instrument from guitar to violin
1 add instrument 139 metal audio with a distortion guitar [and drums] metal audio with a distortion guitar add drums to the piece
2 delete instrument 133 an eerie tense instrumental featuring electronic drums [and synth keyboard] an eerie tense instrumental featuring electronic drums remove the synth keyboard
3 change genre 134 a recording of a solo electric guitar playing [blues] licks a recording of a solo electric guitar playing [rocks] licks change the genre from blues to rock
4 change mood 100 a recording featuring electric bass with an [upbeat] vibe a recording featuring electric bass with an [melancholic] vibe turn upbeat mood into melancholic mood
5 change rhythm 69 a live ukulele performance featuring [fast] strumming and emotional melodies a live ukulele performance featuring [slow] strumming and emotional melodies change fast rhythm into slow one
6 change background 95 female voices in unison with [acoustic] guitar female voices in unison with [electric] guitar switch acoustic guitar to electric guitar
7 change melody 121 this instrumental song features a [relaxing] melody this instrumental song features a [cheerful] melody change relaxing melody into cheerful melody
8 extract instrument 111 a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar] a reggae rhythm recording with bongos extract bongos from the recording
9 random 67 / / /
Table 4: Information of ZoME-Bench dataset
B.2 Annotation Process
We rebuild our caption from captions for Musiccaps offered by  (Agostinelli et al., 2023). With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type "change melody" as examples):
"There is a description of a Piece of music ,Please judge whether the description has information of melody.If not , just answer "Flase" ,else change its melody properly into the opposite one ,just change the adjective and don’t replace any instrument! "blended_word" is [origin melody, changed melody] "emphasize" is [changed melody] "blended_word" and "emphasize" are tuples.
Question:(A mellow, passionate melody from a noisy electric guitar) Answer:("original_prompt": "A mellow, [passionate] melody from a noisy electric guitar", "editing_prompt": "A mellow, [soft] melody from a noisy electric guitar", "blended_word": ["passionate melody","soft melody"], "emphasize" :["soft melody"]) Question:(A recording of solo harp music with a dreamy, relaxing melody.)
Answer:("original_prompt": "A recording of solo harp music with a dreamy, [relaxing] melody.", "editing_prompt": "A recording of solo harp music with a dreamy, [nervous] melody.", "blended_word": ["relaxing melody","nervous melody"], "emphasize" :["nervous melody"])
Question:("A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
Answer:("original_prompt": "A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", "editing_prompt": "A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", "blended_word": ["harmonized flute melody","passionate flute melody"]), "emphasize" :["passionate flute melody"])
now we have Question({origin caption})
Answer(?)"
In the same way , instructions were appended by prompt as follows(take type "change melody" as examples):
"There is two description of different pieces of music divided by &,Please describe the difference you need to give me the results in the following format: Question: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo & this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
Answer: change relaxing melody into cheerful melody
Question:this song features acapella harmonies with a [high pitched] melody complemented by both high pitched female whistle tones and male low pitch tones & this song features acapella harmonies with a [smooth] melody complemented by both high pitched female whistle tones and male low pitch tones
Answer:turn a high pitched melody into smooth melody
Question: a traditional and hopeful song with a harmonizing throaty male vocal and [dissonant] background melody from strings albeit presented in low quality & a traditional and hopeful song with a harmonizing throaty male vocal and [harmonic] background melody from strings albeit presented in low quality
Answer: change dissonant melody into harmonic melody  Question: [’original prompt’] & [’editing prompt’]
Answer(?)"
Through this method accmopnied with rounds of manual review ,we guarantee the quality of this Benchmark.
B.3 Data Format
Taking the first piece as example , we express our data in json format with six keys
"$000$": {
"editing_prompt": "a live recording of ambient acoustic
[violin] music",
"original_prompt": "a live recording of ambient acoustic
[guitar] music",
"blended_word": "(\"guitar\", \"violin\")",
"emphasize": "(\"violin\")",
"audio_path": "wavs/MusicCaps_-4SYC2YgzL8.wav",
"editing_type_id": "0",
"editing_instruction": "change the instrument from guitar
to violin"
"Editing_prompt" is the edited caption while "Original_prompt" is the origin one . "Blended_word" is the subject to be edited."Emphasize" is the word which should be emphasized. "editing_instruction" is the description of editing.Something else that needs attention is in the editing type "delete instrument" , we create another key "neg_prompt" , which helps to reduce the possibility of deleted instruments still appearing.
Appendix C Implementation Details
For our evaluation, we utilize the pre-trained AudioLDM 2-Music model (Liu et al., 2023b). Our assessment employs a comprehensive set of metrics, namely CLAP, LPAPS, Structure Distance, and FAD. These metrics are calculated using the CLAP models available in the AudioLDM_eval package, which is accessible at https://github.com/haoheliu/audioldm_eval. In line with the methodology described by  Manor & Michaeli (2024), we apply a forward guidance of 3 and a reverse guidance scale of 12 for DDPM inversion. For the DDIM inversion, the guidance scale is set to 5, while for SDEdit, we employ a guidance scale of 12. The forward guidance of DIC is 1 while the reverse scale is 5. We chose these values by exploration about different guidance scale, as discussed in Appendix D. We conduct all experiments in NVIDIA 4090. For our evaluation, we have selected the public pre-trained AudioLDM 2-Music model (Liu et al., 2023b). To ensure a thorough and multidimensional assessment, we measure performance using a suite of metrics that includes CLAP, LPAPS, Structure Distance, and FAD, and conduct on an NVIDIA 4090. The computation of these metrics is facilitated by the CLAP models provided within the AudioLDM_eval package, which is publicly available at https://github.com/haoheliu/audioldm_eval.
Our methodology is aligned with the protocol established by  Manor & Michaeli (2024), where we have adopted a forward guidance scale of 3 and a reverse guidance scale of 12 for DDPM inversion. In contrast, the DDIM inversion employs a guidance scale of 5, and SDEdit utilizes a guidance scale of 12. For Disentangled Inversion Control, we have determined the forward guidance to be 1 and the reverse scale to be 5. These specific guidance scale values are selected after extensive experimental exploration, the details of which are discussed in Appendix D.
C.1 Metrics
There are details about five metrics to evaluate the performance of our novel Disentangled Inversion Control framework: (1) Struture Distance (Ju et al., 2023): Leveraging self-similarity of audio features to measure the structure distance between the source and edited audio. (2) CLAP Score (Elizalde et al., ): This criterion evaluates the degree to which the output conforms to the specified target prompt. (3) LPAPS (Iashin & Rahtu, 2021; Paissan et al., 2023): An audio adaptation of the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), this measure evaluates the consistency of the edited audio with the original audio. (4) MSE (Mean Squared Error): This distance metric quantifies the variance between the source and the edited audio. (5) FAD (Fréchet Audio Distance) (Kilgour et al., 2018): Analogous to the FID used in image analysis, this metric calculates the distance between two distributions of audio signals.
Appendix D Quantitative Results
For the lack of systematic experiments which determines the optimal combination of guidance scales for achieving the best editing performance, and analysis on how these guidance scales affect final consequence in both reconstruction and editing, we conduct this experiment to find the best scales.
Guidance Scale Structure Background Preservation CLIP Similariy
[…]
1 1 8.56 0.12 1.17 3.25 0.51
1 2.5 11.97 0.15 2.49 4.54 0.56
1 5 15.99 0.17 4.22 6.07 0.60
1 7.5 15.99 0.17 4.22 6.07 0.59
2.5 1 22.80 0.20 6.39 8.65 0.30
2.5 2.5 14.24 0.16 2.50 5.40 0.46
2.5 5 14.46 0.16 3.31 5.49 0.53
2.5 7.5 15.51 0.17 3.94 5.89 0.53
5 1 29.94 0.24 9.81 11.36 0.20
5 2.5 29.16 0.24 9.11 11.07 0.22
5 5 22.15 0.20 5.59 8.40 0.36
5 7.5 17.57 0.18 5.57 6.67 0.48
7.5 1 31.41 0.25 10.62 11.92 0.20
7.5 2.5 31.05 0.25 10.14 11.78 0.20
7.5 5 29.20 0.24 9.32 11.08 0.24
7.5 7.5 24.16 0.22 7.33 9.17 0.34
Table 5: Ablation Studies on Different Guidance Scale
Appendix E Potential Negative Societal Impacts
MEDIC may also lead to potential negative societal impacts that are worthy of consideration. If the data sample of the training model is not diverse enough or biased, the AI-generated music may be overly biased toward one style or element, limiting the diversity of the music and causing discrimination.MEDIC could be used to create fake audio content, such as faking someone’s voice or creating fake musical compositions, posing the risk of fraud and impersonation.Hopefully , all these issues could be taken into consideration when taking the model for real use to avoid ethical issues.
Appendix F Limitations
In spite of remarkable outcome of our method , but due to the limitation of the generation model we used , we are incapable of instigating a profound change.
Though our method is faster than est-optimization and end-to-end diffusion models,but due to the numerous steps it requires (T=200), the duration of computing distance is quite long.Thus , we will implement a more powerful text-to-music generation model to support better editing, while trying to use consistency model or flow matching model to achieve high quality and fast music generation in future work
Appendix G Qualitative Results
For each type in ZoME-Bench , We provide samples to observe the capability of MEDIC intuitively.
G.1 Change Instrument
In the following figures , we show the capability of MEDIC to change the instrument.Here we edit the ground truth music piece with original prompt "a live recording of ambient acoustic [guitar] music" and editing prompt "a live recording of ambient acoustic [violin] music". The difference of instruments can be observed in the Mel-spectrum
source: a live recording of ambient acoustic [guitar] music
edit: a live recording of ambient acoustic [violin] music
Figure 5: Editing Type 0 :Change Instrument
G.2 Add Instrument
In the following figures , we show the capability of MEDIC to add more instrument.Here we edit the ground truth music piece with original prompt "a heavy metal instructional audio with a distortion guitar" and editing prompt "a heavy metal instructional audio with a distortion guitar [and drums]". The appearance of new instrument can be observed in the Mel-spectrum which present drum sound of high frequency.
source: a heavy metal instructional audio with a distortion guitar
edit: a heavy metal instructional audio with a distortion guitar [and drums]
Figure 6: Editing Type 1 :Add Instrument
G.3 Delete Instrument
In the following figures , we show the capability of MEDIC to delete instrument.Here we edit the ground truth music piece with original prompt "a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood" and editing prompt "a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood". The vanishing of instrument can be observed in the Mel-spectrum
source: a lively ska instrumental featuring keyboard trumpets bass [and percussion] with a groovy mood
edit: a lively ska instrumental featuring keyboard trumpets and bass with a groovy mood
Figure 7: Editing Type 2 :Delete Instrument
G.4 Change Genre
In the following figures , we show the capability of MEDIC to change genre of a music piece. Here we edit the ground truth music piece with original prompt "a recording of a solo electric guitar playing [blues] licks" and editing prompt "a recording of a solo electric guitar playing [rock] licks". The obvious difference of genre can be observed in the Mel-spectrum
source: a recording of a solo electric guitar playing [blues] licks
edit: a recording of a solo electric guitar playing [rock] licks
Figure 8: Editing Type 3 :Change Genre
G.5 Change Mood
Mood is an important attribute of a music. In the following figures ,we show the capability of MEDIC to change mood of a music piece. Here we edit the ground truth music piece with original prompt "a recording of [aggressive] electronic and video game music with synthesizer arrangements"and editing prompt "a recording of [peaceful] electronic and video game music with synthesizer arrangements". The change of Mood can be observed in the Mel-spectrum
source: a recording of [aggressive] electronic and video game music with synthesizer arrangements
edit: a recording of [peaceful] electronic and video game music with synthesizer arrangements
Figure 9: Editing Type 4 :Change Mood
G.6 Change Rhythm
Rhythm represents the speed of the music. In the following figures ,we show the capability of MEDIC to change Rhythm of a music piece. Here we edit the ground truth music piece with original prompt "a [slow] tempo ukelele tuning recording with static"and editing prompt "a [fast] tempo ukelele tuning recording with static". The change of Rhythm can be observed in the Mel-spectrum.The edited Mel-spectrum is much intensive.
source: a [slow] tempo ukelele tuning recording with static
edit: a [fast] tempo ukelele tuning recording with static
Figure 10: Editing Type 5 :Change Rhythm
G.7 Change Background
In the following figures ,we show the capability of MEDIC to change background of instrument of a music piece. Here we edit the ground truth music piece with original prompt "an amateur ukulele recording with a [medium to uptempo] pace"and editing prompt "an amateur ukulele recording with a [steady and rhythmic] pace". The change of instrument background can be observed in the Mel-spectrum.
source: an amateur ukulele recording with a [medium to uptempo] pace"and editing prompt
edit: an amateur ukulele recording with a [steady and rhythmic] pace
Figure 11: Editing Type 6 :Change Background
G.8 Change Melody
In the following figures ,we show the capability of MEDIC to change melody of a music piece. Here we edit the ground truth music piece with original prompt "this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo"and editing prompt "this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo". The change of Melody can be observed in the Mel-spectrum
source: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
Figure 12: Editing Type 7 :Change Melody
G.9 Extract Instrument
In the following figures ,we show the capability of MEDIC to extract one certain instrument of a music piece. Here we edit the ground truth music piece with original prompt "a reggae rhythm recording with bongos [djembe congas acoustic drums and electric guitar]"and editing prompt "a reggae rhythm recording with bongos". The change of instruments can be observed in the Mel-spectrum. Much sounds vanished.
source: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
edit: this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
Figure 13: Editing Type 8 :Extract Instrument
Appendix H safeguards
In the processing of the data and models involved in this study, we fully considered the potential risks. We ensure that all data sources are rigorously screened and vetted,and the model we used is absolutely trained from the safe dataset to minimize security risks of being misused.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We accurately reflect the paper’s contributions and scope in the abstract and section 1 introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in section F limitations
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide theory assumptions and proofs in section 2 preliminaries and analyses
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: we will open-source our codes(contains model or their access method) and the benchmark we constructed, which is enough to reproduce our experiments.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: we specify all the training and test details in section 5 Experiments and section C Implementation Details
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We only use one random seed.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide sufficient information on the computer resources in section 5.1 Implementation Details
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Positive impacts is in section 6 Conclusion.Negative impacts in section E Potential Negative Societal Impacts
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We take actions for safeguards , which is described in section H Safeguards
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Creators or original owners of assets, used in the paper, are properly credited and the license and terms of use explicitly mentioned and are properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our assets is well documented and the document is provided
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work do not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work do not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.